A galvanometer has a coil resistance of Rgalv = 13.9 ?. It is to be made into a voltmeter with a full-scale voltage equal to V = 325.0 V. If the galvanometer deflects full scale for a current of Igalv = 245.5 mA, what size additional resistor should be used?
Answer #1

R = (V / I_galv) - R_galv = (325 / 245.5) kohm - 13.9 ohm = 1.31 kohm